OpenStudy (anonymous):

Find \(\color{red}{x}\). \(\color{green}{y^2 + 2y + 1 = 0}\)

OpenStudy (solomonzelman):

I am not god at math, as you have seen many times before, but maybe I can try .... \(\normalsize\color{black}{y^2+2x+1=0}\) is the equation for a parabola opening to the right.\(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{x=y^2+2x+1}\)\(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{x=(y+1)^2}\)\(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{x=1(y+1)^2+0}\)\(\LARGE\color{white}{ \rm │ }\) All I can say is x in terms of y, x=(y+1)²

OpenStudy (anonymous):

What??

OpenStudy (solomonzelman):

nvm.... I knew it will be bad

OpenStudy (anonymous):

Have I typed the question right or haven't you seen the question with your eyes?

OpenStudy (anonymous):

the letter in red is x

hartnn (hartnn):

i was going to say same thing :P 

OpenStudy (anonymous):

@sauravshakya has found it.. Good...

OpenStudy (solomonzelman):

odd

OpenStudy (anonymous):

@hartnn sorry, saurav was somewhat faster than you in finding that... :)

OpenStudy (anonymous):

the question should be solve for x

OpenStudy (anonymous):

It is not a question @No.name

OpenStudy (anonymous):

The real question is coming now.. :)

OpenStudy (solomonzelman):

bye

OpenStudy (anonymous):

Bye?? Why??? Where you are going??? :) You are on OS already and will be. :)

OpenStudy (yanasidlinskiy):

Wow!!! Awesome question!

OpenStudy (anonymous):

Yeah I know like you.. :) Thank you thank you.. :P

OpenStudy (yanasidlinskiy):

Pshshhsh....I'm not awsomexD!

OpenStudy (anonymous):

Today you know what is common in we both??

OpenStudy (yanasidlinskiy):

I don't wanna know:)

OpenStudy (anonymous):

Okay, you are not but my joke is.. :)

OpenStudy (anonymous):

\[\Huge Expand 1\]

OpenStudy (anonymous):

Expand 1??

OpenStudy (anonymous):

Are you alright @No.name ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Does not seem to be.. :)